1. If a rod travels with a speed v 0.8 c along its length, how much does it shrink? 2. Assuming that the rest radius of earth is 6,400 km and its orbital speed about the sun is 30 km, how much does earths diameter appear to be shortened to an observer on the sun, due to earths orbital motion? 3. A rod of length Lo moves with speed v along the horizontal direction. The rod makes and angle 0o with respect to the x axis. a. Determine the length of the rod as measured by a stationary observer. b. Determine the angle θ the rod makes with the x axis.
Answer #1

=0.6xレ 1.66 0-6ourectton 2.= 12-79 918106m OL = (12-8-12-79 9999 )入106m 0,064-m